# Liam Welsh

### v2

@@ 1 Introduction
Excess emissions are one major cause of man-made climate change. These emissions consequently are causing more severe weather events, such as forest fires, hurricanes, and floods, and displacing millions of people each year. In an attempt to limit the impacts, more governments worldwide (both at the federal level and regional level) are pricing carbon in an attempt to incentivize emissions reductions and transition to greener technologies. There are over sixty carbon pricing systems worldwide that attempt to control carbon emissions, such as carbon taxes and emission markets. Despite this, only 4% of these systems are valuing carbon emissions at a rate that would meet the $2^{\circ}$ C warming limit set by the Paris Agreement (Santikarn et al.,, 2021). This is because there is no unified global price of carbon (International Monetary Fund,, 2022). In an attempt to minimize this disparity, many international treaties and agreements have been made limit the impacts of climate change, including the Montreal Protocol (1987), Kyoto Protocol (1997), the Paris Agreement (2015), and the Glasgow Climate Pact (2021). These agreements all aim to limit excess GHGs, provide financial incentives to implement climate friendly legislation for developing nations, and increase the development of green and renewable technologies. To better track emissions within the financial sector, Kenyon et al., (2022, 2023) introduces the carbon equivalence principle (CEP), which argues that the addition of a term sheet structure for financial products of carbon emissions that are a result of the asset. CEP structured products, they argue, allow for increased carbon visibility and creation of net-zero portfolios, and is found to drive further sustainability efforts.
@@
One method meant to financially incentivize firms to limit emissions in Canada is the implementation of a new GHG OC market (Government of Canada, 2022b, ; Government of Canada, 2022a, ). This market is structured similar to both renewable energy certificate (REC) and carbon cap-&\&-trade (C&\&T) markets, of which there is extensive literature. Seminal work by Seifert et al., (2008) characterize firm behaviour in a C&\&T as the solution to an optimal control problem in a single period setting, which is then extended to a multi-period model in Hitzemann and Uhrig-Homburg, (2018). Carmona et al., (2009) investigates optimal behaviour in carbon markets while simultaneously shedding light on some of the issues within these markets, and in Carmona et al., (2022) the authors analyses the influence that carbon taxes have on energy producers in the mean-field limit. Aïd and Biagini, (2023) solve the Stackelberg game between the regulator, who wishes to reduce overall carbon emissions, and emitting firms, wherein the regulator can dynamically assign emission allowances. In REC markets, regulators provide certificates to firms who produce electricity through renewable means. These markets have also been well studied, including by Amundsen et al., (2006) who describes price volatility models for REC markets. Shrivats and Jaimungal, (2020) solves for a firm’s optimal controls in a solar REC market, which they then extend to a mean-field setting in Shrivats et al., (2022) with endogenized REC pricing, and finally to a principal agent mean-field game in Firoozi et al., (2021). The influence that emissions have on investors and companies has also been well studied. Cartellier et al., (2023) thoroughly and mathematically demonstrates why corporations employ greenwashing techniques, and the influence that environmentally conscience investors have on a firm’s decision to employ such techniques. Greenwashing is a deceptive practice in which corporations exaggerate their green practices and environmental investments. Greenwashing practices are now under significant scrutiny due to international and societal pressures (Sengupta,, 2024). Aldy and Halem, (2024) provides a comprehensive overview of OCs through a policy and regulation perspective, investigating the role OCs have as a compliance device, a voluntary credit for unregulated firms, and an incentive for climate conscious behaviour.
@@
The majority of this prior research utilizes classical techniques, and there is little research on employing modern statistical machine learning, such as reinforcement learning (RL), to model these complex markets. This is burgeoning area due to both the advancements in learning theory techniques and the increasing severity of the impact that man-made climate change is having on our planet. Ahmadi et al., (2024) use a deep RL technique to solve for the optimal trade-off between the mean-variance and the GHG footprint trade-off of a GHG emission adverse portfolio. Barreau et al., (2024) proposes machine learning methodologies to accurately estimate the direct and indirect emissions of firms using public data, allowing quantitative backed scrutiny to be applied to major polluters and firms that under-report the extent of their environmental impact. Noh and Swidinsky, (2025) employ deep RL techniques for geological carbon sequestration to determine the optimal carbon injection rate for storage and to control and minimise geological noise (e.g. seismic activity) while injecting. From the extant literature, the contemporaneous work of Cialenco and Ludkovski, (2025) shares the most similarities to our paper and the prior work by Welsh and Jaimungal, (2024). There, the authors develop a finite-agent ground water market model in which agents optimize their water production with respect to their water rights, consumptions, and trading, and can realize further rewards by producing goods (e.g. crops) using the water. The authors fully characterize and study the resulting Nash equilibrium in the single-period setting. However, due to the computational intractability, they propose using machine learning approaches to characterize the multi-period Nash equilibrium as a future direction.
@@
In this paper, we present a framework for estimating the finite-player Nash equilibrium in GHG OC markets using a deep $Q$-learning technique called Nash-DQN (Casgrain et al.,, 2022) and study the resulting equilibrium. The approach we use is computationally efficient, and through experiments and simulations we demonstrate the viability of the methodology and the financial importance for firms to actively participate in these markets. The remainder of the paper is organized as follows. In Section 2, we provide background on the Canadian GHG market structure that forms the basis of our market model. and present the Nash-DQN methodology in Section 2.2.1. The mathematical GHG OC market model and resulting RL implementation is provided in Section 3 and Section 3.1, respectively. Experimental results are presented in Section 4, wherein we demonstrate the versatility of the methodology. Closing thoughts and potential directions for future research are provided in Section 5.
@@ 2.1 Market Structure
Canadian emissions markets exist both provincially and federally. Provincial governments are allowed to regulate their own GHG emissions framework and market under the assumption that it meets the minimum federal pricing standards and GHG reduction targets (Sadikman et al.,, 2022). In 2022, the Canadian federal government announced a new GHG OC system at the federal level with the goal of expanding project opportunities and streamlining the process (Government of Canada, 2022a, ). In this updated framework, there are compliance markets and voluntary markets that firms may participate in. Firms may be regulated or unregulated, where a regulated firm is given an emissions limit over some period. Regulated firms participate in the compliance market, and if they exceed their limit are penalized proportionally to their excess emissions. To reduce this penalty, firms may submit OCs to the regulating body which they have either generated through project investment or have purchased from some other entity that generated OCs. By submitting OCs to the regulating body, firms reduce their penalty. Any unused OCs may be carried over into future years, up to an OC’s expiry date. OCs generated in the voluntary market may not be used for compliance purposes, due to a less rigid regulatory procedures in obtaining generated OCs. Unregulated firms that choose to participate in the voluntary market (and are not compelled to generate OCs to offset excess emissions nor do they face any potential regulatory penalty) do so to amplify their environmental and social goals, or they may participate due to pressure from environmentally conscious investors. De Angelis et al., (2023) demonstrates the impact that environmentally driven investors can have in curbing a firm’s emissions, and further propose raising the cost of capital for more polluting firms to encourage emission reductions.
@@
The new Canadian federal market began accepting projects for registration in January of 2024, and as of September 2025 there are 32 registered projects and a total of 5,000 OCs issued (Government of Canada,, 2024). Provincial GHG OC markets have long been established prior to this, including in British Columbia (BC) and Québec. In the BC market, 9.3 million Mt of $\text{CO}_{2}\text{e}$ were offset from the public sector between 2010 and 2023 (Government of British Columbia,, 2024), while in Québec’s 1.8 million OCs were issued between January 2014 and July 2024 (Gouvernement du Québec,, 2024). Further, the Québec market structure only allows 8% of a firms emissions to be offset via OCs, hence incentivising general emission reductions (Gouvernement du Québec,, 2024). Provinces who do not have GHG markets or have markets that do not meet federal GHG pricing standards fall under the federal market.
Prior work by Welsh and Jaimungal, (2024) characterizes the optimal behaviour in a single-player model and the mixed-strategy Nash equilibrium in a two-player game by employing classical stochastic control and game theoretic techniques coupled with finite-difference approximations. In this work, we estimate the Nash equilibrium for a finite number of regulated firms participating in the compliance market using deep learning techniques. We assume that these regulated firms all have the ability to invest in and generate their own OCs, or they can trade OCs.
@@ 2.2 Reinforcement Learning and Nash-DQN
Obtaining finite-player Nash equilibria is, in general, an NP-complete problem (Daskalakis et al.,, 2009) — i.e., it cannot be solved efficiently unless $P=NP$. Our approach is therefore to approximate such equilibria using RL techniques. RL (see Sutton, (2018)) combines control theory with machine learning in order to approximate and learn optimal strategies for various problems. In such problems, an agent interacts with an environment — without explicitly knowing its evolution, i.e., using the empirical measure — in an attempt to learn the optimal policy by observing instances of states, actions, and rewards. In short, an agent observes an environmental state, chooses to take some action, and then receives a reward depending on the state, action, and (potentially) the next environmental state.
@@
where $\pi_{i}$ is agent-$i$’s policy, ${\boldsymbol{\pi}}_{-i}:=\{\pi_{1},\dots,\pi_{i-1},\allowbreak\pi_{i+1}\,\dots,\pi_{N}\}$ is the collection of other players’ policies, and $\gamma_{t,i}\in(0,1]$ is a discount factor for agent-$i$,. In the sequel, we use the subscript i-i to denote the set ${\mathfrak{N}}\setminus\{i\}$. An agent’s $Q$-value is defined as the total (discounted) reward of an agent taking the specific action $a$ in the state $\theta$ and following the optimal policy for the remainder of the game. $Q$-learning (see Watkins et al., (1989); Watkins and Dayan, (1992)) is an extremely versatile off-policy learning method for an MDP in which the learned values converge to the $Q$-values. This is accomplished with the Bellman equation. The Bellman equation equates the value of the current state to the sum of the immediate reward and the (discounted) future value, and may be expressed in an operator form. Due to the contraction mapping of the Bellman operator, convergence to the optimal action-state value function is guaranteed in $Q$-learning under mild assumptions (Watkins and Dayan,, 1992). In a multi-agent setting, each agent’s $Q$-value is (typically) dependent on all agents’ policies. For the remainder of this work, we discuss and assume a multi-agent RL setting.
@@
holds for all admissible policies $\pi_{i}:\mathcal{S}\to\mathcal{A}$, for all $i\in{\mathfrak{N}}$, and for all $\theta\in\mathcal{S}$. That is, if any agent deviates from their Nash policy while the other agents follow their Nash policy, the deviating agent is unable improve their performance. We denote $\Pi$ to be the set of admissible policies. For stationary strategies, every stochastic finite-player game has at least one Nash equilibrium in stationary strategies (Fink,, 1964). These equilbria are, however, often difficult to compute and computationally intractable. We utilise RL techniques to efficiently approximate these Nash equilibria.
@@
$Q_{i}(\theta;a_{i},\mathbf{a}_{-i}):=\mathbf{\mathfrak{r}}_{i}(\theta;a_{i},\mathbf{a}_{-i})+\gamma_{i}\;\mathbb{E}_{\theta^{\prime}\sim p(\cdot|\theta,\mathbf{a})}\big[V_{i}(\theta^{\prime})\big],$ (4)
where $V_{i}$ denotes the $i$-th agent’s value function. The notation $\theta^{\prime}\sim p(\cdot|\theta,\mathbf{a})$ indicates that the new state $\theta^{\prime}$ has distribution given by $p(\cdot|\theta,\mathbf{a})$. The $Q$-function may be interpreted as the reward agent-$i$ receives by taking the fixed action $a_{i}$, while others take fixed action $\mathbf{a}_{-i}$, when in state $\theta$, and subsequently taking optimal actions once the new state reveals itself. Collecting the $Q$-function across agents into a vector, as well as the rewards and the value function, we re-write (4) as
@@
for all $i\in{\mathfrak{N}}$, and $f_{i}$ denotes the $i$-th component of $\mathbf{f}$ (see Casgrain et al., (2022) for a formal definition) — we assume here that there is a unique such point. The Nash operator maps a vector of functions that accept a vector of actions to the vector of values at the Nash-equilibria. Using the Nash operator, we can relate the $Q$-functions and the value functions through the Bellman equation (3) as
@@
This equation is the Nash-Bellman equation (Casgrain et al.,, 2022).
@@ 2.2.1 Nash-DQN
To estimate the finite-player Nash equilibria for a GHG OC market, we use the Nash-DQN algorithm of Casgrain et al., (2022), wherein the authors develop a computationally efficient deep $Q$-learning methodology to efficiently estimate the Nash equilibria of stochastic games. The authors accomplish this by combining the iLQG framework from Todorov and Li, (2005) and Gu et al., (2016) with the Nash $Q$-learning algorithm (NashQ) for general sum stochastic games of Hu and Wellman, (2003). Casgrain et al., (2022) employ local linear-quadratic approximations and additional concavity assumptions to efficiently estimate the Nash equilibrium. We provide an overview of the Nash-DQN methodology. In our market setting, discussed in detail in Section 3, each agent has a two-dimensional action space, and hence we accordingly define vectors and matrices with this in mind.
@@
where $\alpha$ represents the parameters of the approximations (e.g., the weights and biases of a NN). This decomposition should be viewed as the defining equation for the advantage function (see Schulman et al., (2015) for generalized advantage function estimation). The advantage function may be seen as the (negative) improvement that the optimal action is in state $\theta$ compared with taking the specific action $\mathbf{a}$ followed by optimal actions afterwards. Using an advantage function in learning, in addition to the value function, mimics the actor-critic approach. In Nash-DQN, each agent’s advantage function is approximated as being locally linear-quadratic with respect to the agents’ actions with coefficients given as outputs of DNNs — by locally linear-quadratic, we mean that the coefficients of the quadratic form in $\mathbf{a}$ vary with state $\theta$. Specifically, in our setting where agents select actions corresponding to both a rate of trading and a probability of generating, the advantage functions, for all $i\in{\mathfrak{N}}$, admit the quadratic-form representation
@@
can be constructed using the output of a single DNN. Given the restriction that $P_{11,i}^{\alpha}(\theta)$ is required to be positive-definite, Casgrain et al., (2022) employ a Cholesky decomposition and instead model the lower triangular matrix, which we mimic here and hence we estimate $L_{i}^{\alpha}(\theta)\in\mathds{R}^{2\times 2}$ and compute $P_{11,i}=L_{i}^{\alpha}(\theta)\left(L_{i}^{\alpha}(\theta)\right)^{\intercal}.$

### v3

@@ 1 Introduction
Excess emissions are one major cause of man-made climate change. These emissions consequently are causing more severe weather events, such as forest fires, hurricanes, and floods, and displacing millions of people each year. In an attempt to limit the impacts, more governments worldwide (both at the federal level and regional level) are pricing carbon in an attempt to incentivize emissions reductions and transition to greener technologies. There are over sixty carbon pricing systems worldwide that attempt to control carbon emissions, such as carbon taxes and emission markets. Despite this, only 4% of these systems are valuing carbon emissions at a rate that would meet the $2^{\circ}$ C warming limit set by the Paris Agreement (Santikarn et al., 2021). This is because there is no unified global price of carbon (International Monetary Fund, 2022). In an attempt to minimize this disparity, many international treaties and agreements have been made to limit the impacts of climate change, including the Montreal Protocol (1987), Kyoto Protocol (1997), the Paris Agreement (2015), and the Glasgow Climate Pact (2021). These agreements all aim to limit excess greenhouse gasses (GHGs), provide financial incentives to implement climate friendly legislation for developing nations, and increase the development of green and renewable technologies. To better track emissions within the financial sector, Kenyon et al. (2022, 2023) introduces the carbon equivalence principle (CEP), which argues that the addition of a term sheet structure for financial products of carbon emissions that are a result of the asset. CEP structured products, they argue, allow for increased carbon visibility and creation of net-zero portfolios, and is found to drive further sustainability efforts.
@@
One method meant to financially incentivize firms to limit emissions in Canada is the implementation of a new GHG offset credit (OC) market (Government of Canada, 2022b, a). This market is structured similar to both renewable energy certificate (REC) and carbon cap-&\&-trade (C&\&T) markets, of which there is extensive literature. Seminal work by Seifert et al. (2008) characterize firm behaviour in a C&\&T as the solution to an optimal control problem in a single period setting, which is then extended to a multi-period model in Hitzemann and Uhrig-Homburg (2018). Carmona et al. (2009) investigates optimal behaviour in carbon markets while simultaneously shedding light on some of the issues within these markets, and in Carmona et al. (2022) the authors analyses the influence that carbon taxes have on energy producers in the mean-field limit. Aïd and Biagini (2023) solve the Stackelberg game between the regulator, who wishes to reduce overall carbon emissions, and emitting firms, wherein the regulator can dynamically assign emission allowances. In REC markets, regulators provide certificates to firms who produce electricity through renewable means. These markets have also been well studied, including by Amundsen et al. (2006) who describes price volatility models for REC markets. Shrivats and Jaimungal (2020) solves for a firm’s optimal controls in a solar REC market, which they then extend to a mean-field setting in Shrivats et al. (2022) with endogenized REC pricing, and finally to a principal agent mean-field game in Firoozi et al. (2021). The influence that emissions have on investors and companies has also been well studied. Cartellier et al. (2025) thoroughly and mathematically demonstrates why corporations employ greenwashing techniques, and the influence that environmentally conscience investors have on a firm’s decision to employ such techniques. Greenwashing is a deceptive practice in which corporations exaggerate their green practices and environmental investments. Greenwashing practices are now under significant scrutiny due to international and societal pressures (Sengupta, 2024). Aldy and Halem (2024) provides a comprehensive overview of OCs through a policy and regulation perspective, investigating the role OCs have as a compliance device, a voluntary credit for unregulated firms, and an incentive for climate conscious behaviour.
@@
The majority of this prior research utilizes classical techniques, and there is little research on employing modern statistical machine learning, such as reinforcement learning (RL), to model these complex markets. This is burgeoning area due to both the advancements in learning theory techniques and the increasing severity of the impact that man-made climate change is having on our planet. Ahmadi et al. (2024) use a deep RL technique to solve for the optimal trade-off between the mean-variance and the GHG footprint trade-off of a GHG emission adverse portfolio. Barreau et al. (2025) proposes machine learning methodologies to accurately estimate the direct and indirect emissions of firms using public data, allowing quantitative backed scrutiny to be applied to major polluters and firms that under-report the extent of their environmental impact. Noh and Swidinsky (2025) employ deep RL techniques for geological carbon sequestration to determine the optimal carbon injection rate for storage and to control and minimise geological noise (e.g. seismic activity) while injecting. From the extant literature, the contemporaneous work of Cialenco and Ludkovski (2025) shares the most similarities to our paper and the prior work by Welsh and Jaimungal (2024). There, the authors develop a finite-agent ground water market model in which agents optimize their water production with respect to their water rights, consumptions, and trading, and can realize further rewards by producing goods (e.g. crops) using the water. The authors fully characterize and study the resulting Nash equilibrium in the single-period setting. However, due to the computational intractability, they propose using machine learning approaches to characterize the multi-period Nash equilibrium as a future direction.
@@
In this paper, we present a framework for estimating the finite-player Nash equilibrium in GHG OC markets using a deep $Q$-learning technique called Nash-DQN (Casgrain et al., 2022) and study the resulting equilibrium. The approach we use is computationally efficient, and through experiments and simulations we demonstrate the viability of the methodology and the financial importance for firms to actively participate in these markets. The remainder of the paper is organized as follows. In Section 2, we provide background on the Canadian GHG market structure that forms the basis of our market model. and present the Nash-DQN methodology in Section 2.2.1. The mathematical GHG OC market model and resulting RL implementation is provided in Section 3 and Section 3.1, respectively. Experimental results are presented in Section 4, wherein we demonstrate the versatility of the methodology. Closing thoughts and potential directions for future research are provided in Section 5.
@@ 2.1 Market Structure
Canadian emissions markets exist both provincially and federally. Provincial governments are allowed to regulate their own GHG emissions framework and market under the assumption that it meets the minimum federal pricing standards and GHG reduction targets (Sadikman et al., 2022). In 2022, the Canadian federal government announced a new GHG OC system at the federal level with the goal of expanding project opportunities and streamlining the process (Government of Canada, 2022a). In this updated framework, there are compliance markets and voluntary markets that firms may participate in. Firms may be regulated or unregulated, where a regulated firm is given an emissions limit over some period. Regulated firms participate in the compliance market, and if they exceed their limit are penalized proportionally to their excess emissions. To reduce this penalty, firms may submit OCs to the regulating body which they have either generated through project investment or have purchased from some other entity that generated OCs. By submitting OCs to the regulating body, firms reduce their penalty. Any unused OCs may be carried over into future years, up to an OC’s expiry date. OCs generated in the voluntary market may not be used for compliance purposes, due to a less rigid regulatory procedures in obtaining generated OCs. Unregulated firms that choose to participate in the voluntary market (and are not compelled to generate OCs to offset excess emissions nor do they face any potential regulatory penalty) do so to amplify their environmental and social goals, or they may participate due to pressure from environmentally conscious investors. De Angelis et al. (2023) demonstrates the impact that environmentally driven investors can have in curbing a firm’s emissions, and further propose raising the cost of capital for more polluting firms to encourage emission reductions.
@@
The new Canadian federal market began accepting projects for registration in January of 2024, and as of September 2025 there are 32 registered projects and a total of 5,000 OCs issued (Government of Canada, 2024). Provincial GHG OC markets have long been established prior to this, including in British Columbia (BC) and Québec. In the BC market, 9.3 million Mt of $\text{CO}_{2}\text{e}$ were offset from the public sector between 2010 and 2023 (Government of British Columbia, 2024), while in Québec’s 1.8 million OCs were issued between January 2014 and July 2024 (Gouvernement du Québec, 2024). Further, the Québec market structure only allows 8% of a firms emissions to be offset via OCs, hence incentivising general emission reductions (Gouvernement du Québec, 2024). Provinces who do not have GHG markets or have markets that do not meet federal GHG pricing standards fall under the federal market.
Prior work by Welsh and Jaimungal (2024) characterizes the optimal behaviour in a single-player model and the mixed-strategy Nash equilibrium in a two-player game by employing classical stochastic control and game theoretic techniques coupled with finite-difference approximations. In this work, we estimate the Nash equilibrium for a finite number of regulated firms participating in the compliance market using deep learning techniques. We assume that these regulated firms all have the ability to invest in and generate their own OCs, or they can trade OCs.
@@ 2.2 Reinforcement Learning and Nash-DQN
Obtaining finite-player Nash equilibria is, in general, an NP-complete problem (Daskalakis et al., 2009) — i.e., it cannot be solved efficiently unless $P=NP$. Our approach is therefore to approximate such equilibria using RL techniques. RL (see Sutton (2018)) combines control theory with machine learning in order to approximate and learn optimal strategies for various problems. In such problems, an agent interacts with an environment — without explicitly knowing its evolution, i.e., using the empirical measure — in an attempt to learn the optimal policy by observing instances of states, actions, and rewards. In short, an agent observes an environmental state, chooses to take some action, and then receives a reward depending on the state, action, and (potentially) the next environmental state.
@@
where $\pi_{i}$ is agent-$i$’s policy, ${\boldsymbol{\pi}}_{-i}:=\{\pi_{1},\dots,\pi_{i-1},\allowbreak\pi_{i+1}\,\dots,\pi_{N}\}$ is the collection of other players’ policies, and $\gamma_{t,i}\in(0,1]$ is a discount factor for agent-$i$,. In the sequel, we use the subscript i-i to denote the set ${\mathfrak{N}}\setminus\{i\}$. An agent’s $Q$-value is defined as the total (discounted) reward of an agent taking the specific action $a$ in the state $\theta$ and following the optimal policy for the remainder of the game. $Q$-learning (see Watkins and others (1989); Watkins and Dayan (1992)) is an extremely versatile off-policy learning method for an MDP in which the learned values converge to the $Q$-values. This is accomplished with the Bellman equation. The Bellman equation equates the value of the current state to the sum of the immediate reward and the (discounted) future value, and may be expressed in an operator form. Due to the contraction mapping of the Bellman operator, convergence to the optimal action-state value function is guaranteed in $Q$-learning under mild assumptions (Watkins and Dayan, 1992). In a multi-agent setting, each agent’s $Q$-value is (typically) dependent on all agents’ policies. For the remainder of this work, we discuss and assume a multi-agent RL setting.
@@
holds for all admissible policies $\pi_{i}:\mathcal{S}\to\mathcal{A}$, for all $i\in{\mathfrak{N}}$, and for all $\theta\in\mathcal{S}$. That is, if any agent deviates from their Nash policy while the other agents follow their Nash policy, the deviating agent is unable improve their performance. We denote $\Pi$ to be the set of admissible policies. For stationary strategies, every stochastic finite-player game has at least one Nash equilibrium in stationary strategies (Fink, 1964). These equilbria are, however, often difficult to compute and computationally intractable. We utilise RL techniques to efficiently approximate these Nash equilibria. We define the value function $V_{i}$ of the agent such that $V_{i}(\theta)=R_{i}(\theta;\,\pi_{i}^{*},\,\boldsymbol{\pi}_{-i}^{*})$, where the value function represents the sum of the current reward and (discounted) expected future rewards under all agents abiding by their equilibrium policies.
@@
$Q_{i}(\theta;a_{i},\mathbf{a}_{-i}):=\mathbf{\mathfrak{r}}_{i}(\theta;a_{i},\mathbf{a}_{-i})+\gamma_{i}\;\mathbb{E}_{\theta^{\prime}\sim p(\cdot|\theta,\mathbf{a})}\big[V_{i}(\theta^{\prime})\big].$ (4)
The notation $\theta^{\prime}\sim p(\cdot|\theta,\mathbf{a})$ indicates that the new state $\theta^{\prime}$ has distribution given by $p(\cdot|\theta,\mathbf{a})$. The $Q$-function may be interpreted as the reward agent-$i$ receives by taking the fixed action $a_{i}$, while others take fixed action $\mathbf{a}_{-i}$, when in state $\theta$, and subsequently taking optimal actions once the new state reveals itself. Collecting the $Q$-function across agents into a vector, as well as the rewards and the value function, we re-write (4) as
@@
for all $i\in{\mathfrak{N}}$, and $f_{i}$ denotes the $i$-th component of $\mathbf{f}$ (see Casgrain et al. (2022) for a formal definition) — we assume here that there is a unique such point. The Nash operator maps a vector of functions that accept a vector of actions to the vector of values at the Nash-equilibria. Using the Nash operator, we can relate the $Q$-functions and the value functions through the Bellman equation (3) as
@@
This equation is the Nash-Bellman equation (Casgrain et al., 2022).
@@ 2.2.1 Nash-DQN
To estimate the finite-player Nash equilibria for a GHG OC market, we use the Nash-DQN algorithm of Casgrain et al. (2022), wherein the authors develop a computationally efficient deep $Q$-learning methodology to efficiently estimate the Nash equilibria of stochastic games. The authors accomplish this by combining the iLQG framework from Todorov and Li (2005) and Gu et al. (2016) with the Nash $Q$-learning algorithm (NashQ) for general sum stochastic games of Hu and Wellman (2003). Casgrain et al. (2022) employ local linear-quadratic approximations and additional concavity assumptions to efficiently estimate the Nash equilibrium. We provide an overview of the Nash-DQN methodology. In our market setting, discussed in detail in Section 3, each agent has a two-dimensional action space, and hence we accordingly define vectors and matrices with this in mind.
@@
where $\alpha$ represents the parameters of the approximations (e.g., the weights and biases of a NN). This decomposition should be viewed as the defining equation for the advantage function (see Schulman et al. (2015) for generalized advantage function estimation). The advantage function may be seen as the (negative) improvement that the optimal action is in state $\theta$ compared with taking the specific action $\mathbf{a}$ followed by optimal actions afterwards. That is, the advantage function can be viewed as providing a measure of action quality, separating the value of taking a specific action in a given state from the optimal action when in that state. By decomposing the Q-function into value and advantage components, the model can explicitly distinguish between the intrinsic value of a state and the relative benefit of selecting a particular action. Omitting this decomposition, standard Q-learning learns only the combined state–action value, making it more difficult to isolate the contribution of individual actions. Work by Tang et al. (2023) demonstrates that directly learning value and advantage functions is advantageous to traditional Q-learning. Using advantage functions in Q-learning have also been shown to improve learning stability, and can further help reduce variance in the policy gradient estimation. Using an advantage function in learning, in addition to the value function, mimics the actor-critic approach. In Nash-DQN, each agent’s advantage function is approximated as being locally linear-quadratic with respect to the agents’ actions with coefficients given as outputs of DNNs — by locally linear-quadratic, we mean that the coefficients of the quadratic form in $\mathbf{a}$ vary with state $\theta$. Specifically, in our setting where agents select actions corresponding to both a rate of trading and a probability of generating, the advantage functions, for all $i\in{\mathfrak{N}}$, admit the quadratic-form representation
@@
can be constructed using the output of a single DNN. Given the restriction that $P_{11,i}^{\alpha}(\theta)$ is required to be positive-definite, Casgrain et al. (2022) employ a Cholesky decomposition and instead model the lower triangular matrix, which we mimic here and hence we estimate $L_{i}^{\alpha}(\theta)\in\mathds{R}^{2\times 2}$ and compute $P_{11,i}=L_{i}^{\alpha}(\theta)\left(L_{i}^{\alpha}(\theta)\right)^{\intercal}.$